OpenStudy (anonymous):

What is the Stefan Boltzmann Radiation constant for ? Want a brief description.

OpenStudy (anonymous):

The Stefan–Boltzmann constant (also Stefan's constant), a physical constant denoted by the Greek letter σ, is the constant of proportionality in the Stefan–Boltzmann law: the total energy radiated per unit surface area of a black body in unit time is proportional to the fourth power of the thermodynamic temperature. The value of the Stefan–Boltzmann constant is given in SI units by σ = 5.670373(21)×10−8 W m−2 K−4.[1]

OpenStudy (anonymous):

@Parvathysubhash where is this constant actually used?

OpenStudy (anonymous):

In cgs units the Stefan-Boltzmann constant is: . In US customary units the Stefan-Boltzmann constant is [2]: . The value of the Stefan–Boltzmann constant is derivable as well as experimentally determinable; see Stefan–Boltzmann law for details. It can be defined in terms of the Boltzmann constant as:

OpenStudy (anonymous):

The Stefan-Boltzmann law: R=σ*T^4; P-power; P=R*S;